OpenStudy (anonymous):

6-7u=3u^2 step 1: -3u^2-7u+6=0 step 2: -(3u^2+7u-6)=0 step 3: (3u-2)(u+3)=0 (Answer in the book) My Question: In step 3: Why didn't we use 1 and 6 (for 6's Greatest Common Factor) which will result in 7 (linear term). 2 and 3 either give 5 or 1. Please clarify I'm confused?

OpenStudy (luigi0210):

Because you're also distributing that \(3u\)

OpenStudy (anonymous):

Oh ok. How could I have missed that. Thank you!

OpenStudy (anonymous):

So I just use a different GCF it's not random?

OpenStudy (mathstudent55):

\(3u^2 + 7u - 6 \) \( = 3u^2 -2u + 9u - 6\) \(=u(3u - 2) +3(3u - 2) \) \(=(3u - 2)(u + 3) \)